OpenStudy (anonymous):

If the following fraction is reduced, what will be the exponent on the p? -5p^5q^4/8p^2q^2 a) 5 b) 4 c) 3 d) 2

OpenStudy (anonymous):

3, exponents of the same base that are being divided by each other are subtracted

OpenStudy (anonymous):

as long as they are positive that is, if a number has a negative exponent, generally you move that number to the opposite end of the fraction. ex: 3^-2 / 6^-3 = 6^3 / 3^2

OpenStudy (anonymous):

also keep in mind that this can only be done if there are no +/- in the equation. its against mathematical rules. ex: 3x^2 7y/6x^7 = 3(7y) / 6x^5 however, 3x^2 + 7y/6x^7 DOES NOT = 3 + 7y / 6x^5

OpenStudy (anonymous):

glad i could help